OpenStudy (cry_baby):

The Howlter family wants to enclose their circular garden with a fence. (a) To the nearest foot, what is the minimum length of fencing they need to buy? (b) If the fence costs $2.75 per foot, what will be the total cost? (Answer both parts/sentence) (the fence is pictured below.)

satellite73 (satellite73):

the circuference of a circle with radius \(r\) is \(2\pi r\)

satellite73 (satellite73):

compute \[2\times \pi \times 50.5\] with a calculator, then round to the nearest whole number

satellite73 (satellite73):

i guess i should say "round up"